OpenStudy (anonymous):

CAN SOMEONE ANSWER THIS QUESTIOM Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. - tan2x + sec2x = 1

OpenStudy (raden):

u meant - tan^2 x + sec^2 x = 1 ??

OpenStudy (anonymous):

yes sorry

OpenStudy (raden):

ok, use the identity : sec^2 x = tan^2 x + 1 now, apply it to left side.

OpenStudy (anonymous):

i got it down to [-sin^2 x +1]/cos^2x but i dont know where to go from there

OpenStudy (raden):

ohh, different by my way. remember the identity : sin^2 x + cos^2 x = 1, so 1-sin^2 x = cos^2 x

OpenStudy (anonymous):

can you just show me step by step how to do it? im so confused

OpenStudy (raden):

i think u almost finishing, just the last step : convert [-sin^2 x +1] = cos^2 x so, [-sin^2 x +1]/cos^2 x = cos^2 x/cos^2 x = 1 right ?

OpenStudy (anonymous):

yes, thank you so much!

OpenStudy (raden):

welcome